# Collective steering: Tracer-informed dynamics

### v1

@@ Abstract
We consider control and inference problems where control protocols and internal dynamics are informed by two types of constraints. Our data consist of i) statistics on the ensemble and ii) trajectories or final disposition of selected tracer particles embedded in the flow. Our aim is i’) to specify a control protocol to realize a flow that meets such constraints or ii’) to recover the internal dynamics that are consistent with such a data set. We analyze these problems in the setting of linear flows and Gaussian distributions. The control cost is taken to be a suitable action integral constrained by either the trajectories of tracer particles or their terminal placements.
@@ 1 Introduction
The theme of this work is the control problem of steering a collection of particles/agents that obey a universal time-varying feedback law; the law represents a common control protocol broadcast to all agents in the collective. In this paper, we are interested in both control and inference problems. The data of these problems concern, on the one hand, the collective in the form of marginal distribution constraints and, on the other, the locations or trajectories of specific tracer particles embedded in the flow. Whether we consider inference or control, these tracer particles reveal internal dynamics or constrain a sought-after control protocol.
@@
The practical motivation for this class of problems stems from a range of applications in engineering and physics. To mention a few, this is the case in particle systems steered by externally regulated control potentials. Similarly, in swarm robotics, exogenous control inputs are used to steer the swarm into particular configurations. On the flip side, landmarks in the field of view of a collective may serve as tracer particles’ constraints to guide the motion of the ensemble overflying a known terrain. Indeed, also in fluids, tracer particles may reveal the internal dynamics of the flow.
@@
The present work aims to develop this circle of ideas in the most basic setting of linear flows and Gaussian distributions. Accordingly, we consider linear time-varying feedback control laws. Control costs penalize both the kinetic energy of the ensemble as well as the control gain. The former is physically motivated and underlies the theory of optimal mass transport [1, 2, 3, 4]. The latter quantifies the complexity of control implementation as was eloquently explained by R.W. Brockett [5].
@@
In linear state feedback, when the feedback gain is regulated and represents the control authority, the dynamics are bilinear. Due to this fact, controllability issues are subtle and often overlooked. For instance, while it was widely accepted that in this basic setting, the Liouville equation is controllable under standard conditions [6], a proof of this fact was not available until recently [7] (see also, [8]). At the heart of the technical issue is that solutions to optimal control problems for linear systems may not always be expressible in linear feedback form due to topological obstructions [8, Section IV]. Thus, a main contribution of the present work, besides raising attention to this type of problems where tracer particles may inform control and inference, is to establish the existence of solutions to such problems under suitable regularizations of the cost.
@@
Below, we introduce notation and discuss the general setting in Section 2. In Section 3, we introduce and solve the first of our problems: the problem of collective steering with the information on the running distribution of the collective in conjunction with the endpoint positions of tracers. In Section 4, we formulate and solve a dual problem entailing information on the endpoint distributions of the collective and trajectories of the tracers. Lastly, we present an academic example for each problem in Section 5.
@@ 2 Problem setting
Throughout, we consider a collection of particles in transport over the time window $t\in[0,1]$ that are subject to the linear dynamics
@@
${\rm Sym}^{+}(n):=\{\Sigma\in\mathbb{R}^{n\times n}:\Sigma\succ 0,\Sigma=\Sigma^{\top}\},$
i.e., the space of $n\times n$ positive-definite and symmetric matrices. If ${\rm GL}(n)$ denotes the real general linear group of degree $n$, i.e., the multiplicative group of $n\times n$ invertible matrices, then the flow of (1) takes the form
@@
$\dot{\Phi}_{t}=K_{t}\Phi_{t},\ \ \Phi_{0}=I.$ (3)
As is well-known, the space of Gaussian distributions is invariant under the action of linear flows, i.e., flows of the form (2), via the pushforward of measures. Then, with direct computation, one can verify that $X_{t}\sim\mathcal{N}_{n}(0,\Phi_{t}\Sigma_{0}\Phi_{t}^{\top})$.
@@
where $\Sigma_{t}\in\mathrm{Sym}^{+}(n)$. We will only be interested in the cases where $\mathcal{T}_{1}=\{0,1\}$ or $\mathcal{T}_{1}=[0,1]$. In the latter case, we shall always assume that the curve $\Sigma_{t\in[0,1]}$ is continuously differentiable.
The second type of constraints, ii), can be succinctly explained as follows. Let $y_{t}^{(i)}$ be the position of the $i$-th tracer particle at the time instant $t\in\mathcal{T}_{2}\subseteq[0,1]$ for $i\in\{1,\ldots m\}$ with $m<n$, and let
@@
Once again, we will only be interested in the cases where $\mathcal{T}_{2}=\{0,1\}$ or $\mathcal{T}_{2}=[0,1]$, and in the latter case, we shall always assume that the curve $Y_{t\in[0,1]}$ is continuously differentiable and of full column rank.
To wrap up this preliminary section, we now introduce the cost functionals we deal with. The first control cost we consider is the integral of the (expected) kinetic energy
$\frac{1}{2}\int_{0}^{1}\mathbb{E}[\|\dot{X}_{t}\|^{2}]\,\mathrm{d}t,$ (6)
where $\mathbb{E}[\cdot]$ denotes the expectation operator (recall that $X_{0}\sim\mathcal{N}(0,\Sigma_{0})$, i.e., $X_{0}$ is a Gaussian random variable). The functional (6) is the typical cost in optimal mass transport and quantifies the control energy, see, e.g., [2, 4]. Substituting (1) and (2) into (6) while utilizing the properties of the expectation operator and the fact that $\mathbb{E}[X_{0}X_{0}^{\top}]=\Sigma_{0}$, we obtain that the functional (6) can be rewritten as
$\int_{0}^{1}{\rm tr}(K_{t}\Phi_{t}\Sigma_{0}\Phi_{t}^{\top}K_{t}^{\top})\,\mathrm{d}t.$ (7)
Notice that, in terms of $K_{t}$ and $\Phi_{t}$, the cost functional (7) is not jointly convex, and so establishing the existence of solutions of optimal control problems with the cost (7) is non-trivial. Indeed, for some terminal specifications the minimizing (7) leads to a singular state transition $\Phi_{t}$ and, consequently, a non-integrable $K_{t}$ [8, Example 1]. The second cost functional we consider is
$\int_{0}^{1}\mathbb{E}[\|\nabla u(X_{t})\|^{2}]\,\mathrm{d}t,$ (8)
@@
$\int_{0}^{1}{\rm tr}(K_{t}K_{t}^{\top})\,\mathrm{d}t.$ (9)
Note that unlike the functional (7), the cost functional (9) is convex in $K_{t}$ and is independent of $\Phi_{t}$. It turns out that utilizing the functional (9) as a regularization term for the cost functional (7) is instrumental in establishing the existence of solutions to one of the problems we consider.
3 Internal Dynamics from Tracer Displacements
In this section, we tackle the first of the problems we propose in this paper. Throughout this section, we assume knowledge about the ensemble of particles in the form of the marginal distributional constraints (4) with $\mathcal{T}_{1}=[0,1]$ and the endpoint conditions on the tracers (5) with $\mathcal{T}_{2}=\{0,1\}$. Our task is to identify the dynamics in (3), effected by a suitable control vector field $u_{t}$ as in (1), that realizes the given data about the flow. We take the functional (7) as our control cost. The problem we investigate can then be formulated as follows.
@@ 3 Internal Dynamics from Tracer Displacements
Given $\Sigma_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=[0,1]$ and $Y_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=\{0,1\}$, find the minimizer of (7), subject to the constraints (3), (4), and (5).
Let us define what shall be viewed as an admissible solution to the previous problem. We say a curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ is admissible for Problem 1 if $K_{t\in[0,1]}$ is Lebesgue-integrable and $\Phi_{t\in[0,1]}$ is absolutely continuous, and are such that the differential constraint (3) is satisfied (whenever $\dot{\Phi}_{t}$ exists) as well as constraints (4) and (5), and the value of the functional (7) is finite. The following proposition provides a more explicit characterization of the class of admissible curves.
@@
second, the constraint (5) is satisfied, third, there exists an almost everywhere unique Lebesgue-integrable curve $\Omega_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ such that
$K_{t}=\frac{1}{2}\dot{\Sigma}_{t}\Sigma_{t}^{-1}+\Omega_{t}\Sigma_{t}^{-1},\ \ \mbox{ with }\ \ \Omega_{t}^{\top}=-\Omega_{t},$ (11)
@@
Let $(K_{t},\Phi_{t})_{t\in[0,1]}$ be an admissible curve for Problem 1. Since $\mathcal{T}_{1}=[0,1]$ and $\Sigma_{t\in\mathcal{T}_{1}}$ is continuously differentiable by assumption, direct differentiation of (4) followed by substitution of (3) shows that $K_{t\in[0,1]}$ must satisfy the Lyapunov equation
@@
for almost all $t\in[0,1]$. On the other hand, since $\Sigma_{t}\in\mathrm{Sym}^{+}(n)$, it follows from basic linear algebra that the constraint (12) is satisfied for all $t\in\mathcal{T}_{1}=[0,1]$ if and only if $K_{t}$ is given by (11), where the Lebesgue-integrable curve $\Omega_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ is almost everywhere uniquely defined. Finally, since $1\in\mathcal{T}_{1}=[0,1]$, the curve $\Phi_{t\in[0,1]}$ must be such that $\Phi_{1}$ satisfies (10). The converse implication follows by direct integration of the differential constraint (3) with $K_{t\in[0,1]}$ defined by (11).
Next, we observe that if $(K_{t},\Phi_{t})_{t\in[0,1]}$ is any admissible curve for Problem 1, the functional (7) is given by
$\int_{0}^{1}{\rm tr}(K_{t}\Sigma_{t}K_{t}^{\top})\,\mathrm{d}t,$
where we have substituted the constraint (4) into the functional (7). The advantage of the latter representation of the functional is that it is convex in the free parameter $\Omega_{t}$ (after another substitution from (11)) and is independent of $\Phi_{t}$. The preceding discussion suggests the reformulation of Problem 1 as follows.
@@
Given $\Sigma_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=[0,1]$ and $Y_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=\{0,1\}$, find the minimizer of
$\int_{0}^{1}{\rm tr}(K_{t}\Sigma_{t}K_{t}^{\top})\,\mathrm{d}t$
@@
Suppose at least one admissible curve exists for Problem 1. Then, there exists an absolute minimizer $(K_{t},\Phi_{t})_{t\in[0,1]}$ for Problem 1.
The theorem implies that Problem 1 has an absolute minimizer, which, however, may not be unique. The proof, due to its technical nature, is omitted. Briefly, it consists of verifying that the conditions of [9, Theorem 11.4.vi] hold. Thus, in light of Theorem 2, we can now meaningfully seek a minimizer by solving the necessary conditions of optimality obtained from the classical theory of optimal control [10] as formulated next.
@@
Let $(K_{t},\Phi_{t})_{t\in[0,1]}$ be an absolute minimizer of Problem 1, or equivalently Problem 1. Then,
$\dot{\Phi}_{t}$ $=K_{t}\Phi_{t},$ (13a)
with $\Phi_{0}=I$ and $K_{t}$ is given by (11) and
$\Omega_{t}\Sigma_{t}+\Sigma_{t}\Omega_{t}$ $=$
$\frac{1}{2}(\Phi_{t}P_{t}^{\top}+\dot{\Sigma}_{t})\Sigma_{t}-\frac{1}{2}\Sigma_{t}(P_{t}\Phi_{t}^{\top}+\dot{\Sigma}_{t}),$ (13b)
where $\text{where }\hskip 20.0pt\dot{P_{t}}$ $=-{K_{t}}^{\top}P_{t}.$ (13c)
@@
then (13a) and (13c) are subject to the mixed boundary conditions $\Phi_{1}\in\mathcal{S}$ and ${P_{1}}\perp T_{\Phi_{1}}\mathcal{S}$, respectively.
@@
$\mathcal{H}$ $=\frac{1}{2}{\rm tr}\big{(}K_{t}\Sigma_{t}K_{t}^{\top}+P_{t}^{\top}K_{t}\Phi_{t}\big{)},$
@@
for $K_{t}$ satisfying (11) and where $P_{t}$ is an $n\times n$ Lagrange multiplier (costate) matrix enforcing (3). The first order necessary conditions in $P_{t}$ and $\Phi_{t}$ are such that $\dot{\Phi}_{t}=\partial\mathcal{H}/\partial P_{t}$ and $\dot{P}_{t}=-\partial\mathcal{H}/\partial\Phi_{t}$, respectively. These two conditions yield Equations (13a) and (13c), in that order.
Using the form (11), we rewrite the previous Hamiltonian as
$\mathcal{H}$ $=\frac{1}{2}{\rm tr}\big{(}\big{(}\frac{1}{2}\dot{\Sigma}_{t}+\Omega_{t}\big{)}\Sigma_{t}^{-1}\big{(}\frac{1}{2}\dot{\Sigma}_{t}+\Omega_{t}\big{)}^{\top}\big{)}$
$+{\rm tr}\big{(}\big{(}P_{t}^{\top}\big{(}\frac{1}{2}\dot{\Sigma}_{t}+\Omega_{t}\big{)}\Sigma_{t}^{-1}\Phi_{t}\big{)},$
@@
bearing in mind that $\Omega_{t}$ is skew-symmetric. In this case, Pontryagin’s maximum principle is such that $\partial\mathcal{H}/\partial\Omega_{t}=0$. Using that and the fact that
$\frac{\partial}{\partial\Omega_{t}}{\rm tr}\big{(}\Omega_{t}S_{t}\big{)}=\frac{1}{2}\big{(}S_{t}-S_{t}^{\top}\big{)},$
for $S_{t}\in\mathbb{R}^{n\times n}$, the Lyapunov equation (13b) can be obtained.
@@
The terminal condition (10) along with tracers’ condition (5) define the surface $\mathcal{S}$ in (13d). Thus, it is necessary that $\Phi_{1}\in\mathcal{S}$. According to [10, Section 5.1], any admissible variation for the optimal $\Phi_{1}$, $\delta\Phi_{1}$, has to be tangent to $\mathcal{S}$, i.e., $\delta\Phi_{1}\in T_{\Phi_{1}}\mathcal{S}$. We note,
${\rm tr}({P_{1}}^{\top}\delta\Phi_{1})=0,$
@@
for any admissible variation $\delta\Phi_{1}$, cf. [10, Equation 5.1-18], which necessitates that ${P_{1}}\perp T_{\Phi_{1}}\mathcal{S}$.
@@
Any solution to Problem 1 must satisfy the necessary conditions of optimality stated in Proposition 3, constituting a nonlinear two-point boundary value problem. A solution to this problem, which is guaranteed to exist as a consequence of Theorem 2, provides a candidate flow that reconciles the data of Problem 1 with the particle dynamics (1) using a suitable linear feedback law. Numerical results, in addition to visualization of the flow, are presented in Section 5 to illustrate the nature of the solutions to this problem.
4 Internal Dynamics from Tracer Trajectories

### v2

@@ Abstract
We consider control and inference problems where control protocols and internal dynamics are informed by two types of constraints. Our data consist of i) statistics on the ensemble and ii) trajectories or final dispositions of selected tracer particles embedded in the flow. Our aim is i’) to specify a control protocol, realizing a flow that meets such constraints or ii’) to recover the internal dynamics that are consistent with such a data set. We analyze these problems in the setting of linear flows and Gaussian distributions. The control cost is taken to be a suitable action integral constrained by either the trajectories of tracer particles or their terminal placements.
@@ 1 Introduction
The theme of this work is the problem of steering a collection of particles/agents that obey a universal time-varying feedback law; the law represents a common control protocol broadcast to all agents in the collective. In this paper, we are interested in both control and inference problems. The data for these problems concern, on one hand, the collective in the form of marginal distribution constraints, and on the other, the locations or trajectories of specific tracer particles, namely tagged agents in the flow. Whether we consider inference or control, the tracer particles are to reveal internal dynamics or constrain a sought-after control protocol.
@@
The practical motivation for this class of problems stems from a range of applications in engineering, physics, and biology. To mention a few, this is the case in particle systems steered by externally regulated control potentials [1, 2]. Similarly, in swarm robotics, exogenous control inputs are used to steer the swarm to desired configurations [3, 4]. On the flip side, landmarks in the field of view of a collective may serve as tracer particles’ constraints to guide the motion of the ensemble traversing a known terrain. Indeed, the use of tracers in medical imaging should not be overlooked [5, 6]. Lastly, tracer particles can also be used in fluids to reveal the internal dynamics of the flow.
@@
The present work aims to develop this circle of ideas in the most basic setting of linear flows and Gaussian distributions. In this, we consider linear time-varying feedback control laws. Control costs penalize both the kinetic energy of the ensemble as well as the control gain. The former is physically motivated and underlies the theory of $L_{2}$-optimal mass transport [7, 8, 9, 10, 11, 12]. The latter quantifies the complexity of control implementation as was proposed and eloquently explained by R.W. Brockett [13].
The average kinetic energy has been a typical cost to transportation problems due to its convexity. Notably, the interpolation of measures given in [14], and known as McCann’s interpolation, coincides with the solution to the $L_{2}$-optimal transport problem. Moreover, in the context of control systems, [15] and [16] considered minimizing the control energy to steer an ensemble that obeys linear dynamics between the endpoint distributions, thereby solving an $L_{2}$-optimal transport problem with prior dynamics. Further, [17] extended the problem to tracking a sequence of distributions based on output measurements, instead of assuming full-state knowledge. Here, we also refer to the review article [18] and the references within.
In the present work, given the data on the collective and tracers, we shall be concerned with identifying a suitable pair of a feedback gain and a state transition matrix that reconciles with the data. While seeking feedback gain is ubiquitous in control problems, identifying associated state transitions is novel and, in fact, essential in our exposition. In that, the purpose of obtaining the state-transitions is to recover the internal degrees of freedom which can not be inferred from solving the $L_{2}$-optimal transport problem; under the assumptions of linear flows and Gaussian distributions, the geometric study [12] (see also [19] and [11]) explains that the $L_{2}$-optimal transport is a geodesic problem on the general linear group. The geodesics constituting the solution to optimal transport solely utilize the horizontal distribution, which is identified with gradient vector fields. Therefore, the problem is reduced to a geodesic problem on the space of Gaussian densities, where geodesics remarkably coincide with those of McCann’s interpolation. This is to say, solving the $L_{2}$-optimal transport problem can only provide information about the underlying flow based on the symmetric part of the state transitions. As such, we are motivated here to obtain state transitions in full, utilizing tracer-based knowledge.
@@
In the case at hand, when the feedback gain and the state transitions are decision variables, the dynamics turn out to be bilinear. Due to this fact, controllability issues are subtle and often overlooked. For instance, while it was widely accepted that, in this basic setting, the Liouville equation is controllable under standard conditions [20], a proof of this fact was not available until recently [21] (see also [22]). At the heart of such a technical issue is the fact that solutions to optimal control problems for linear systems may not always be expressible in linear feedback form due to topological obstructions [22, Section IV]. Thus, besides raising attention to this class of problems where tracer particles may inform control and inference, it is worthwhile to shed light on the delicacy in establishing the existence of solutions to such problems, possibly under suitable regularization of the cost.
@@
Below, we introduce notation and discuss the general setting in Section 2. In Section 3.1, we introduce and solve the first of our problems: the problem of collective steering with the information on the running distribution of the collective in conjunction with the endpoint positions of tracers. In Section 3.2, we formulate and solve a dual problem involving information on the collective’s endpoint distributions and trajectories of the tracers. Before concluding, we present an academic example for each problem in Section 4.
@@ 2 Problem setting
Throughout, we consider a collection of particles in transport over the time window $t\in[0,1]$ that are subject to the linear dynamics:
@@
${\rm Sym}^{+}(n):=\{\Sigma\in\mathbb{R}^{n\times n}~|~\Sigma\succ 0,\Sigma=\Sigma^{\top}\},$
i.e., the space of $n\times n$ positive-definite symmetric matrices. If ${\rm GL}(n)$ denotes the real general linear group of degree $n$, i.e., the multiplicative group of $n\times n$ invertible matrices, then the flow of (1) takes the form
@@
$\dot{\Phi}_{t}=K_{t}\Phi_{t},\quad\mbox{ and }\Phi_{0}=I,$ (3)
for $I$ denoting the identity matrix. As is well-known, the space of Gaussian distributions is invariant under the action of linear flows, i.e., flows of the form (2), via the pushforward of measures. Then, with direct computation, one can verify that $X_{t}\sim\mathcal{N}_{n}(0,\Phi_{t}\Sigma_{0}\Phi_{t}^{\top})$.
@@
where $\Sigma_{t}\in\mathrm{Sym}^{+}(n)$. We will only be interested in the cases where $\mathcal{T}_{1}=\{0,1\}$ or $\mathcal{T}_{1}=[0,1]$. In the latter case, we shall always assume that the curve $(\Sigma_{t})_{t\in[0,1]}$ is continuously differentiable.
The second type of constraints, ii), can be succinctly explained as follows. Let $y_{t}^{(i)}$ be the position of the $i$-th tracer particle at the time instant $t\in\mathcal{T}_{2}\subseteq[0,1]$ for $i\in\{1,\ldots m\}$ and let
@@
Once again, we will only be interested in the cases where $\mathcal{T}_{2}=\{0,1\}$ or $\mathcal{T}_{2}=[0,1]$, and in the latter case, we shall always assume that the curve $(Y_{t})_{t\in[0,1]}$ is continuously differentiable and of full column rank.
We note that in the limiting case $m=n$ and by (5), the matrix $\Phi_{t}=Y_{t}^{-1}Y_{0}$, i.e., the state transitions are fully determined from information on the tracer trajectories. In what follows, we shall assume that $m<n$, and we will be concerned with determining the state transitions that reconcile with given information on the flow via optimization problems. Specifically, we will be concerned with minimizing a cost functional over a to-be-characterized admissible set of pairs $(K_{t},\Phi_{t})$. We shall refer to the global minimum in the admissible set as the absolute minimum, following the terminology of [23].
Before concluding this section, we introduce the cost functionals we will be dealing with. The first control cost we consider is the integral of the (expected) kinetic energy:
$J_{\rm KE}:=\frac{1}{2}\int_{0}^{1}\mathbb{E}[\|\dot{X}_{t}\|^{2}]\,\mathrm{d}t,$ (6)
where $\mathbb{E}[\cdot]$ denotes the expectation operator (recall that $X_{0}\sim\mathcal{N}_{n}(0,\Sigma_{0})$, i.e., $X_{0}$ is a Gaussian random variable). The functional in (6) is a typical cost in optimal mass transport and quantifies the control energy, see, e.g., [9] and [12]. Substituting (1) and (2) into (6) while utilizing the properties of the expectation operator and the fact that $\mathbb{E}[X_{0}X_{0}^{\top}]=\Sigma_{0}$, we obtain that (6) can be rewritten as
$J_{\rm KE}=\frac{1}{2}\int_{0}^{1}{\rm tr}(K_{t}\Phi_{t}\Sigma_{0}\Phi_{t}^{\top}K_{t}^{\top})\,\mathrm{d}t.$ (7)
Notice that, in terms of $K_{t}$ and $\Phi_{t}$, the cost functional in (7) is not jointly convex, and so establishing the existence of solutions of optimal control problems with the cost in (7) is non-trivial. Indeed, for some terminal specifications, minimizing the cost in (7) leads to a singular state transition $\Phi_{t}$ and, consequently, a non-integrable $K_{t}$ [22, Example 1]. The second cost functional we consider is expressed as
$J_{\rm A}:=\frac{1}{2}\int_{0}^{1}\mathbb{E}[\|\nabla u(X_{t})\|^{2}]\,\mathrm{d}t,$ (8)
@@
$J_{\rm A}=\frac{1}{2}\int_{0}^{1}{\rm tr}(K_{t}K_{t}^{\top})\,\mathrm{d}t.$ (9)
Note that, unlike the functional in (7), the cost functional in (9) is convex in $K_{t}$ and is independent of $\Phi_{t}$. It turns out that utilizing the functional in (9) as a regularization term for the cost functional in (7) is instrumental in establishing the existence of solutions to one of the problems we consider.
3 Main Contributions
3.1 Internal Dynamics from Tracer Displacements
Here, we tackle the first of the problems we propose in this paper. Throughout this subsection, we assume knowledge about the ensemble of particles in the form of the marginal distributional constraints (4) with $\mathcal{T}_{1}=[0,1]$ and the endpoint conditions on the tracers (5) with $\mathcal{T}_{2}=\{0,1\}$. Our task is to identify the dynamics in (3), affected by a suitable control vector field $u_{t}$ as in (1), that realizes the given data about the flow. We take the functional in (7) as our control cost. The problem we investigate can then be formulated as follows.
@@ 3.1 Internal Dynamics from Tracer Displacements
Given $(\Sigma_{t})_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=[0,1]$ and $(Y_{t})_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=\{0,1\}$, find
$(K_{t}^{\star},\Phi_{t}^{\star}):=\arg\min_{(K_{t},\Phi_{t})}J_{\rm KE}$
subject to the constraints (3), (4), and (5).
Let us define what shall be viewed as an admissible solution to the previous problem. We say a curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ is admissible for Problem 1 if $(K_{t})_{t\in[0,1]}$ is square integrable and $(\Phi_{t})_{t\in[0,1]}$ is absolutely continuous, and are such that the differential constraint (3) is satisfied (whenever $\dot{\Phi}_{t}$ exists) as well as constraints (4) and (5), and the value of the functional (7) is finite. The following proposition provides a more explicit characterization of the class of admissible curves.
@@
second, the constraint (5) is satisfied, third, there exists an almost everywhere unique square-integrable curve $(\Omega_{t})_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ such that
$K_{t}=\frac{1}{2}\dot{\Sigma}_{t}\Sigma_{t}^{-1}+\Omega_{t}\Sigma_{t}^{-1},\quad\mbox{ with }\Omega_{t}^{\top}=-\Omega_{t},$ (11)
@@
Let $(K_{t},\Phi_{t})_{t\in[0,1]}$ be an admissible curve for Problem 1. Since $\mathcal{T}_{1}=[0,1]$ and $(\Sigma_{t})_{t\in\mathcal{T}_{1}}$ is continuously differentiable by assumption, direct differentiation of (4) followed by substitution of (3) shows that $(K_{t})_{t\in[0,1]}$ must satisfy the Lyapunov equation:
@@
for almost all $t\in[0,1]$. On the other hand, since $\Sigma_{t}\in\mathrm{Sym}^{+}(n)$, it follows from basic linear algebra that the constraint (12) is satisfied for all $t\in\mathcal{T}_{1}=[0,1]$ if and only if $K_{t}$ is given by (11), where the square integrable-curve $(\Omega_{t})_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ is almost everywhere uniquely defined. Finally, since $1\in\mathcal{T}_{1}=[0,1]$, the curve $(\Phi_{t})_{t\in[0,1]}$ must be such that $\Phi_{1}$ satisfies (10). The converse implication follows by direct integration of the differential constraint (3) with $(K_{t})_{t\in[0,1]}$ defined by (11).
Next, we observe that if $(K_{t},\Phi_{t})_{t\in[0,1]}$ is any admissible curve for Problem 1, the functional in (7) is equivalent to
$\frac{1}{2}\int_{0}^{1}{\rm tr}(K_{t}\Sigma_{t}K_{t}^{\top})\,\mathrm{d}t,$
where we have substituted the constraint (4) into (7). The advantage of the latter representation of the functional is that it is convex in the free parameter $\Omega_{t}$ (after another substitution from (11)) and is independent of $\Phi_{t}$. The preceding discussion suggests the reformulation of Problem 1 as follows.
@@
Given $(\Sigma_{t})_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=[0,1]$ and $(Y_{t})_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=\{0,1\}$, find
$(K_{t}^{\star},\Phi_{t}^{\star}):=\arg\min_{(K_{t},\Phi_{t})}\frac{1}{2}\int_{0}^{1}{\rm tr}(K_{t}\Sigma_{t}K_{t}^{\top})\,\mathrm{d}t$
@@
Suppose at least one admissible curve exists for Problem 1. Then, there exists an absolute minimizer $(K_{t}^{\star},\Phi_{t}^{\star})_{t\in[0,1]}$ for Problem 1.
The theorem implies that Problem 1 has an absolute minimizer, which, however, may not be unique. The proof, due to its technical nature, is omitted. Briefly, it consists of verifying that the conditions of [23, Theorem 11.4.vi] hold. Thus, in light of Theorem 2, we can now meaningfully seek a minimizer by solving the necessary conditions of optimality obtained from the classical theory of optimal control [24] as formulated next.
@@
Let $(K_{t}^{\star},\Phi_{t}^{\star})_{t\in[0,1]}$ be an absolute minimizer of Problem 1, or equivalently Problem 1. Then,
$\dot{\Phi}_{t}^{\star}$ $=K_{t}^{\star}\Phi_{t}^{\star},$ (13a)
with $\Phi_{0}^{\star}=I$, $K_{t}^{\star}$ coincides with (11) and $\Omega_{t}$ satisfies
$\begin{split}\Omega_{t}\Sigma_{t}+\Sigma_{t}\Omega_{t}&=\ \big{(}\Phi_{t}^{\star}(P_{t}^{\star})^{\top}+\frac{1}{2}\dot{\Sigma}_{t}\big{)}\Sigma_{t}\\ &-\Sigma_{t}\big{(}P_{t}^{\star}(\Phi_{t}^{\star})^{\top}+\frac{1}{2}\dot{\Sigma}_{t}\big{)},\end{split}$ (13b)
$\hskip-30.0pt\text{where }\hskip 20.0pt\dot{P_{t}}^{\star}$ $=-({K_{t}}^{\star})^{\top}P_{t}^{\star}.$ (13c)
@@
then (13a) and (13c) are subject to the mixed boundary conditions $\Phi_{1}^{\star}\in\mathcal{S}$ and ${P_{1}}\perp T_{\Phi_{1}^{\star}}\mathcal{S}$, respectively.
@@
$\mathcal{H}(K_{t},\Phi_{t},P_{t})$ $=\frac{1}{2}{\rm tr}\big{(}K_{t}\Sigma_{t}K_{t}^{\top})+{\rm tr}(P_{t}^{\top}K_{t}\Phi_{t}\big{)},$
@@
for $K_{t}$ satisfying (11) and where $P_{t}$ is an $n\times n$ Lagrange multiplier (costate) matrix that enforces (3). The first order necessary conditions in $P_{t}$ and $\Phi_{t}$ are such that $\dot{\Phi}_{t}^{\star}=\partial\mathcal{H}/\partial P_{t}$ and $\dot{P}_{t}^{\star}=-\partial\mathcal{H}/\partial\Phi_{t}$, respectively. These two conditions yield Equations (13a) and (13c), in that order.
Using (11), we rewrite the previous Hamiltonian as
$\mathcal{H}(\Omega_{t},\Phi_{t},P_{t})$ $=\frac{1}{2}{\rm tr}\big{(}\big{(}\frac{1}{2}\dot{\Sigma}_{t}+\Omega_{t}\big{)}\Sigma_{t}^{-1}\big{(}\frac{1}{2}\dot{\Sigma}_{t}+\Omega_{t}\big{)}^{\top}\big{)}$
$+{\rm tr}\big{(}P_{t}^{\top}\big{(}\frac{1}{2}\dot{\Sigma}_{t}+\Omega_{t}\big{)}\Sigma_{t}^{-1}\Phi_{t}\big{)},$
@@
bearing in mind that $\Omega_{t}$ is skew-symmetric. In this case, Pontryagin’s maximum principle is such that $\partial\mathcal{H}/\partial\Omega_{t}=0$, from which the Lyapunov equation (13b) can be obtained.
@@
The terminal condition (10) along with tracers’ condition (5) define the surface $\mathcal{S}$ in (13d). Thus, it is necessary that $\Phi_{1}^{\star}\in\mathcal{S}$. According to [24, Section 5.1], any admissible variation for the optimal value $\Phi_{1}^{\star}$, $\delta\Phi_{1}^{\star}$, has to be tangent to $\mathcal{S}$, i.e., $\delta\Phi_{1}^{\star}\in T_{\Phi_{1}^{\star}}\mathcal{S}$. We note,
${\rm tr}(({P_{1}^{\star}})^{\top}\delta\Phi_{1}^{\star})=0,$
@@
for any admissible variation $\delta\Phi_{1}^{\star}$, cf. [24, Equation 5.1-18], which necessitates that ${P_{1}^{\star}}\perp T_{\Phi_{1}^{\star}}\mathcal{S}$.
@@
Any solution to Problem 1 must satisfy the necessary conditions of optimality stated in Proposition 3, constituting a nonlinear two-point boundary value problem. A solution to this problem, which is guaranteed to exist as a consequence of Theorem 2, provides a candidate flow that reconciles the data of Problem 1 with the particle dynamics (1) using a suitable linear feedback law. Numerical results, in addition to visualization of the flow, are presented in Section 4.1 to illustrate the nature of the solutions to this problem.
3.2 Internal Dynamics from Tracer Trajectories